The average angular acceleration of a stone lodged in a rotating car wheel is 2 100 rads . What will the stone’s final angular speed be after 3.0 s, if it starts from rest? Also calculate its angular displacement during this time.



1
Expert's answer
2021-10-13T09:07:39-0400

The angular speed will be


"\omega=\alpha t=100·3=300\text{ rad/s}."

The angular displacement, starting from rest, is


"\phi=\frac{\alpha t^2}2=450\text{ rad}."